# Impact of evanescent scattering modes and finite dispersion in superconducting junctions

### v1

@@ III.4 Cross-channel coupling and evanescent modes
The previous sections dealt with energy dependence due to a single dispersive channel. We now discuss the energy dependence for multi-channel scattering in more detail, where evanescent modes play a role. For simplicity, but without loss of generality, we consider the case where the chemical potential $\mu$ is chosen such that still only the lowest channel is occupied as a planar mode $n_{\mathrm{P}}=1$, i.e. ${\xi_{k=0,n=2}=\mu_{2}>0}$, but now tuned to values where the system is close to occupying the second channel. For $E$ close to zero, we then only have one planar mode (which can propagate forward and backward), and the scattering matrix for electrons (and likewise for holes) has dimensions of two by two (see also blue dashed line in Fig. 3). Consequently, this regime would (according to the state of the art in Sec. II) be described by a simple two-by-two scattering matrix.
But now, with the chemical potential tuned close to the second channel, we immediately see that the scattering matrix inevitably must depend strongly on $E$. Very simply, if we increase the energy $E$ beyond the critical value ${\mu_{\text{crit}}\equiv\xi_{k=0,n=n_{\text{P}}+1}}$ (which for the current example is the minimum of the second channel, i.e. ${\mu_{\mathrm{crit}}=\mu_{2}}$ [^5] Note that we can easily generalize to similar transitions at higher energy, by $\mu_{n_{\text{crit}}}\equiv\xi_{k=0,n=n_{\text{crit}}}$ where $n_{\text{crit}}$ is the index of the channel undergoing the transition., the electron or hole gets two planar channels for its disposal (red dashed line). It stands to reason that the scattering matrix thus changes from two-by-two to four-by-four as a function of $E$, a transition which, in line with the notation adopted in Eq. (6), can be formally depicted as (see also graphical representation in Fig. 3)
@@
Importantly, this observation can be made much more concrete and strongly generalized at the same time. Breaking translation invariance means that the wave function inside the scattering region can no longer be separated into longitudinal and transversal components as in Eq. (8). Hence, solving the usual continuity and differentiability conditions of the scattering problem creates interdependencies between coefficients of different channels.
Figure 4: Dispersion relation of electron bands in the normal conductor. Unoccupied higher planar modes (blue curves) turn into virtually occupied evanescent modes with inverted dispersion below $\mu_{\mathrm{crit}}$ (dashed blue curves). Scattering is energy sensitive if $\mu$ is in the vicinity of one of the minima $\mu_{n}$ (gray transition region).
@@
These real exponentials now satisfy ${\xi_{\pm iK,n}=E}$ for all ${n>n_{P}}$, where $n_{P}$ are the number of planar modes. They can be conveniently represented as inverted dispersion relations (c.f. the dashed blue curves in Figure 4). For a normalizable scattering wave function ansatz, the sign $\pm$ in the above definition has of course to be chosen such that evanescent modes decay far away from the scatterer. However, as we will explicitly show below, for Andreev bound states, our ansatz must include both decaying and diverging modes due to the additional presence of the SN interface.
@@
Consequently, the full scattering problem in the most general case is not described by a finite dimensional matrix (with dimensionality $2n_{P}$) involving only the incoming and outgoing planar modes ${\sim e^{\pm ik_{n}(E)x}}$ (${n\leq n_{P}}$), but is, as a matter of fact, an infinite dimensional matrix describing in addition virtual scattering between planar and evanescent modes, ${\sim e^{\pm K_{n}(E)x}}$ (${n>n_{P}}$), and even between evanescent and evanescent modes. Again, the details of this will be fleshed out in the subsequent section.
@@
Here, we instead continue with the second important conclusion. As indicated above, the transition points where the number of planar channels change are fully dominated by one small energy scale, $\mu_{\text{crit}}$, for the channel ${n_{\text{P}}+1}$ which is evanescent for energies below $\mu_{\text{crit}}$ and planar beyond $\mu_{\text{crit}}$. Consequently, momenta can no longer be approximated to be constant ${k_{n}\left(E\right)\not\approx\mathrm{const.}}$ if ${E\approx\mu_{n}}$. Instead, assuming the dispersion relation given in Eqs. (4) and (5) indicates that momenta ${k_{n}\left(E\right)}$ around the local extremum $\mu_{\text{crit}}$ are highly energy sensitive, following a square root behavior. Note that this is true for energies both above and below $\mu_{\text{crit}}$. The only difference is that the channel ${n_{\text{P}}+1}$ either exhibits a strongly varying imaginary wave vector $\mathrm{i}K_{n_{\text{P}}+1}(E)$ (in the evanescent regime) or a strongly varying real wave vector $k_{n_{\text{P}}+1}(E)$ (planar). In either case, taking the usual differentiability conditions, momentum drops down ${\sim k_{n}\left[E\right]\cdot\psi}$ and enters in first power. Accordingly, it is the momentum $k_{n_{\mathrm{P}}+1}\left(E\right)$ of the channel ${n_{\text{P}}+1}$  and this parameter only  by which energy dependence enters the scattering problem. If scattering is not translational invariant, as discussed above, the dependence of $k_{n_{\mathrm{P}}+1}\left(E\right)$ is projected onto all remaining channels via cross channel interaction, such that the overall energy dependence is dominated by the dispersion of the channel ${n_{\text{P}}+1}$. Thus, we can already predict on this general level, that the behavior of the scattering matrix around that transition is given by the square root law ${\sim\sqrt{|1-E/\mu_{\text{crit}}|}}$.
@@
As a consequence, if the energy is close to the point where additional channels enter the scattering problem, the scattering coefficients smoothly, but strongly, adjust as a function of energy, with the square root power law. That is, there is a region of transition in the vicinity of $\mu_{\text{crit}}$ where the scattering matrix exhibits a strong algebraic energy dependence even below the point where the next higher channel joins the scattering process (c.f. the gray stripe in Figure 4). The full picture, then, is that by approaching $\mu_{\text{crit}}$ from below, evanescent modes gradually become increasingly extended within the conductor arms and eventually turn into planar channels as the next higher mode is populated. In current treatments, it is assumed that evanescent modes exciting the scattering center quickly fade away and thus can safely be neglected. However, as we show below in more detail, for short junctions this assumption is in general valid only for energies sufficiently far away from $\mu_{\text{crit}}$. If the chemical potential is tuned to values close to a channel number transition, especially the lowest evanescent mode (whose spatial extension is diverging here) has to be taken into account, as it essentially pierces the SN-interfaces (c.f. Figure 1). Crucially, this requires a generalization of the interference condition, Eq. (17).
@@
Finally, the above analysis is not only important for the fundamental understanding of weak link physics in the general multi-channel case, but also useful for explicit calculations. As we will show below with the concrete example, the square root behavior can be used as an efficient way to interpolate the scattering matrix at different values of $E$, and thus significantly increase the computational speed.
III.5 Challenges for the generalization of the gap closing argument
@@
As we just demonstrated, when the chemical potential is tuned close to a value where the number of planar channels changes (described by the small energy scale $\mu_{\text{crit}}$), we have a finite energy dependence of the scattering matrix, in spite of negligible dwelling time. It would therefore seem reasonable to expect that the discussion of Sec. III.3 applies here as well, in that the finite energy dependence of the scattering matrix would initially open a gap (i.e., detach $E$ from $\Delta$), but that a correction in $R$ would eventually restore the gap closing. And while ultimately, at least parts of that expectation turn out to be true, the pathway towards a full understanding of the near-gap spectral properties of the ABS in this regime is much more complicated. In this final part of Sec. III, we point out the two main reasons why the generalization of the above gap closing arguments is challenging.
@@
First, there is a simple energy scaling argument indicating that we need to explicitly include evanescent modes into the Beenakker framework. Namely, for the single channel case (see Sec. III.3) we saw that the only relevant energy scale is $\mu$, and that the energy-dependent corrections of the scattering matrix $\sim 1/\mu$ exactly cancelled with the dispersive correction in $R$ (likewise $\sim 1/\mu$). If we were to simply apply the same principle to the multi-channel case, but $\mu$ still tuned to values where there is only one planar channel, we now understand that the planar part of the scattering matrix has corrections that scale with ${1/\mu_{\text{crit}}\gg 1/\mu}$, and therefore, cannot simply be corrected by a mere dispersive correction to the planar Andreev reflection scaling with $1/\mu$. Indeed, as we show below, we need to take into account evanescent modes not only within the scattering matrix describing the central scattering region, but also the evanescent equivalent of Andreev reflection (that is, not only $S$ but also $R$ needs to be extended). Indeed, for the latter, the energy scale $\mu_{\text{crit}}$ emerges again, such that we have at our disposition a process depending on the same, dominant energy scale.
@@
This, however, would not be sufficient, as there is an even more profound challenge. Namely, with Eq. (43), we derived for the single channel case a constraint on the scattering matrix beyond unitarity (but still rooted in probability conservation), under the condition that the finite energy dependence in $S$ originated from a finite dispersion of the conduction electrons. Indeed, it was precisely this constraint that allowed us to show that $E=\Delta$ remained the correct solution at $\phi=0$ [see discussion around Eq. (45)]. It turns out that we can generalize the argument leading to Eq. (43) to the present multichannel case – but this generalized condition will not be helpful. Namely, one can extend the scattering wave function ansatz of Eqs. (29) and (30) to include the dominant evanescent mode in addition to the planar mode, with an evanescent version of a transition and reflection amplitude. However, note that in order for the scattering wave function to be normalizable, we can only include the decaying mode (and not the diverging mode) – an important caveat, as we will see in a moment. Now, proceeding in the same way as in Sec. III.2, one can derive a very similar type of condition as in Eqs. (39) and (42), except that the terms $\sim(R-R^{*})/\mu$ and ${\sim(T-T^{*})/\mu}$ would be replaced by terms proportional to the evanescent reflection and transmission amplitudes, divided by $\mu_{\text{crit}}$. Interestingly, this new constraint would tell a story in between the previously encountered situations. Here, the dip in the probability of the planar mode occupation upon impact of the wave packet at the scattering region is very much real. However, the total occupation probability of the electrons inside the conductor arms is still conserved, because at impact, the evanescent mode is temporarily occupied [c.f. Figure 5b)].
Figure 5: Probability conservation at the time of scattering, depending on the size of the scattering region or on the presence of evanescent modes. A Gaussian wave packet is prepared and impinging upon a scattering center. Unitarity of the $S$-matrix ensures probability conservation before ($t_{1}$) and after ($t_{3}$) the scattering in the conductor arms. However, during the scattering process (around $t_{2}$), properties beyond unitarity become visible, depending on the situation. If the scattering center has non-negligible width [panel a)], part of the probability current flows into the scattering region and is temporarily lost for the conductor arms (leading to a finite energy-dependence of $S$). If we consider instead the case of cross-channel scattering with dispersive evanescent modes [blue real exponentials in panel b)], probability current temporarily flows into these virtually occupied higher channels, even if the scattering region is short (negligible dwelling time ${\tau_{\mathrm{dw}}=1/E_{\mathrm{Th}}\to 0}$). In this latter case, the finite energy-dependence of the planar part of $S$ is compensated through a constraint related to planar-evanescent (and evanescent-planar) scattering processes, cf. Eq. (80) (one of the central results of this work).
@@
Yet, crucially, this constraint would not be a suitable condition to understand the properties of the ABS spectrum near the gap, for the simple reason that the ansatz for the full ABS wave function requires (contrary to the scattering wave function) both decaying as well as diverging parts of the evanescent modes (again, for details see below). The updated probability conservation constraint involving decaying modes however would only provide a relationship between the energy-dependence of the scattering amplitudes of the planar modes and the decaying evanescent modes, and thus provides only incomplete information. As a last resort, one could try to come up with an even more general ansatz for an incident wave packet including both decaying and diverging evanescent modes. But, it is impossible to construct normalizable wave packets in this way.
@@
By carefully generalizing the Beenakker formalism to include evanescent modes below, we will be able to attack this problem from a different angle. In fact, in a sense, we will invert the problem by first providing a perturbative expansion of the generalized Beenakker equation up to first order in $1/\mu_{\text{crit}}$. Then, instead of providing a condition for the scattering matrix such that a gap closing at $\Delta$ is guaranteed, we demand that the gap closing exists, and deduce from that condition a constraint on the generalized scattering matrix (including evanescent modes). Crucially (as indicated above), this constraint cannot follow from probability conservation arguments. Thus, we provide an example where Andreev physics can provide us with fundamental information about the central scattering region, that cannot be obtained otherwise.
@@ IV Beenakker equation including evanescent modes
In order to include evanescent modes into the ansatz of the ABS wave function, we have to find generalizations for both $S$ and $R$ in Eqs. (1) and (10). As already indicated, we need to include in general both decaying and diverging evanescent modes within the unproximitized part of the conductor arms.
Overall, the wave function now needs to have a generalized notion of incoming and outgoing wave amplitudes, $\psi^{\text{in}}$ and $\psi^{\text{out}}$. This distinction is obvious only for planar modes, where one simply distinguishes the propagation direction via the group velocity. For evanescent modes, such a distinction is meaningless, such that we simply choose to define the solutions that decay away from the scattering matrix as outgoing, or “scattered”, evanescent modes (i.e., their amplitudes are added in $\psi^{\text{out}}$), and the ones that diverge as incoming, or “incident”, modes ($\psi^{\text{in}}$), see Fig. 6. We stress that this choice is arbitrary. The resulting wave function ansatz is nonetheless correct, provided that the scattering at the central scattering region, and at the SN interface, are computed consistently with the ensuing asymptotic boundary conditions away from the scattering region.
As a first step, the structure of the scattering matrix as shown in Eq. (6) has to be generalized to include evanescent modes. That is, instead of the submatrices $S^{n,m}$ being limited to indices $n,m$ going from $1$ to $n_{\mathrm{P}}$, where $n_{\mathrm{P}}$ is the number of planar modes, $n,m$ now have (at least in general) no upper bound. For illustration purposes, we explicitly separate the total (infinitely dimensional) scattering matrix into subblocks
@@
As just indicated, since there is an infinite number of higher evanescent modes, the S-matrix now becomes infinitely dimensional. Consequently, for explicit calculations, one has to find an appropriate cutoff for the number of modes, to have convergent low-energy solutions. As we will see in a moment, for realistic problems, it actually often suffices to include only the first evanescent mode, whereas all higher modes can be neglected.
With evanescent modes included into the model, there are real exponentials exiting the scattering center now. Thus, for short junctions where the bulk superconductors are situated really close by [c.f. Figure 1b)], we have to also explicitly solve the Andreev scattering problem at the SN interface with evanescent modes in the N region. Since for $S$, we classified modes decaying from the scattering center as “outgoing” (appearing in $\psi^{\text{out}}$) and diverging modes as “incoming” (appearing in $\psi^{\text{in}}$), we have to apply the same consistent grouping for the boundary conditions at the SN interface as well – now for both electrons and holes. While the wave function ansatz on the N side has both decaying and diverging evanescent components, on the S side, we again have to restrict the ansatz to only decaying modes when receding from the SN interface within the S region (see Fig. 6), to ensure normalizability of the total ABS wave function.
@@
In the state of the art, only planar wave Andreev reflection was included [c.f. Eq. (11)]. Moreover, all coefficients were computed in the limit of a strongly linear dispersion relation (Andreev approximation), such that there was no electron-electron (or hole-hole) backscattering, and the Andreev reflection coefficient $\alpha$ did not depend on the channel index. All of these features must now be included. That is, instead of the form given in the standard framwork by Eq. (11), the total reflection matrix at the SN interface now assumes the form
@@
The matrices $\widehat{\alpha}_{\nu}$ and $\widehat{\beta}_{\nu}$ are diagonal matrices whose size is given by the number of channels times the number of terminals, in analogy to the construction of the matrix ${\text{e}^{-\text{i}\widehat{\phi}}}$ as defined in Eq. (16). The matrix $\widehat{\alpha}_{\nu}$ generalizes Andreev reflection, whereas $\widehat{\beta}_{\nu}$ captures the aforementioned normal backscattering. The coefficients on the diagonal depend on the channel index (but not the terminal index) due to the nonlinearity of each channel’s dispersion. As we will discuss in detail, there is in particular a marked difference in the behaviour of the coefficients for planar- and evanescent modes. As already discussed in Sec. III.3, finite normal backscattering ($\widehat{\beta}\neq 0$) is important especially if it is the planar channels that exhibit strong dispersion. In what follows, we show that if dispersion instead stems from an evanescent mode, corrections to $\widehat{\alpha}$ will play a pivotal role.
@@
$\alpha_{\nu,n}$ $=2\nu\frac{\Delta}{E}\frac{\kappa_{\nu,n}k_{\mathrm{R},n}}{k_{\mathrm{R},n}\zeta_{n}^{-}+\frac{\Delta}{E}\sqrt{1-\frac{E^{2}}{\Delta^{2}}}\Gamma_{n}^{\nu,+}}\ ,$ (62)
@@
$\Gamma_{n}^{\nu,\pm}=$ $\left[\left(k_{\mathrm{I},n}\pm\kappa_{\nu,n}\right)\left(k_{\mathrm{I},n}+\kappa_{-\nu,n}\right)+k_{\mathrm{R},n}^{2}\right]\text{ ,}$
$\zeta_{n}^{\pm}=$ $\left(\kappa_{\mathrm{e},n}\pm\kappa_{\mathrm{h},n}\right)\ .$ (64)
@@
The planar and evanescent momenta ${k_{\mathrm{e},n}}$ and ${k_{\mathrm{h},n}}$ are obtained by considering the dispersion in the N-region [c.f. Eq. (4)] and solve it for the momentum $k_{n}$, where evanescent channels require the replacement ${k_{n}\to-\mathrm{i}K_{n}}$, such that
@@
$K_{\nu,n}$ $=\sqrt{2m\left(\epsilon_{n}-\mu-\nu E\right)}\ .$ (66)

### v2

@@ III.4 Cross-channel coupling and evanescent modes
The previous sections dealt with energy dependence due to a single dispersive channel. We now discuss the energy dependence for multi-channel scattering in more detail, where evanescent modes play a role. For simplicity, but without loss of generality, we consider the case where the chemical potential $\mu$ is chosen such that still only the lowest channel is occupied as a planar mode ${n_{\mathrm{P}}=1}$, i.e., ${\xi_{k=0,n=2}=\mu_{2}>0}$, but now tuned to values where the system is close to occupying the second channel. For $E$ close to zero, we then only have one planar mode (which can propagate forward and backward), and the scattering matrix for electrons (and likewise for holes) has dimensions of two by two (see also blue dashed line in Fig. 4). Consequently, this regime would (according to the state of the art in Sec. II) be described by a simple two-by-two scattering matrix.
But now, with the chemical potential tuned close to the second channel, we immediately see that the scattering matrix inevitably must depend strongly on $E$. Very simply, if we increase the energy $E$ beyond the critical value ${\mu_{\text{crit}}\equiv\xi_{k=0,n=n_{\text{P}}+1}}$ (which for the current example is the minimum of the second channel, i.e., ${\mu_{\mathrm{crit}}=\mu_{2}}$) [^4] Note that we can easily generalize to similar transitions at higher energy, by $\mu_{n_{\text{crit}}}\equiv\xi_{k=0,n=n_{\text{crit}}}$ where $n_{\text{crit}}$ is the index of the channel undergoing the transition., the electron or hole gets two planar channels for its disposal (red dashed line). It stands to reason that the scattering matrix thus changes from two-by-two to four-by-four as a function of $E$, a transition which, in line with the notation adopted in Eq. (6), can be formally depicted as (see also graphical representation in Fig. 4)
@@
Importantly, this observation can be made much more concrete and strongly generalised at the same time. Breaking translation invariance means that the wave function inside the scattering region can no longer be separated into longitudinal and transversal components as in Eq. (8). Hence, solving the usual continuity and differentiability conditions of the scattering problem creates interdependencies between coefficients of different channels.
Figure 5: Dispersion relation of electron bands in the normal conductor. Unoccupied higher planar modes (blue curves) turn into virtually occupied evanescent modes with inverted dispersion below $\mu_{\mathrm{crit}}$ (dashed blue curves). Scattering is energy sensitive if $\mu$ is in the vicinity of one of the minima $\mu_{n}$ (grey transition region).
@@
These real exponentials now satisfy ${\xi_{\pm i\kappa,n}=E}$ for all ${n>n_{P}}$, where $n_{P}$ are the number of planar modes. They can be conveniently represented as inverted dispersion relations (cf. the dashed blue curves in Figure 5). For a normalisable scattering wave function ansatz, the sign $\pm$ in the above definition has of course to be chosen such that evanescent modes decay far away from the scatterer. However, as we will explicitly show below, for Andreev bound states, our ansatz must include both decaying and diverging modes due to the additional presence of the SN interface.
@@
Consequently, the full scattering problem in the most general case is not described by a finite dimensional matrix (with dimensionality $2n_{P}$) involving only the incoming and outgoing planar modes ${\sim e^{\pm ik_{n}(E)x}}$ (${n\leq n_{P}}$), but is, as a matter of fact, an infinite dimensional matrix describing in addition virtual scattering between planar and evanescent modes, ${\sim e^{\pm\kappa_{n}(E)x}}$ (${n>n_{P}}$), and even between evanescent and evanescent modes. Again, the details of this will be fleshed out in the subsequent section.
@@
Here, we instead continue with the second important conclusion. As indicated above, the transition points where the number of planar channels change are fully dominated by one small energy scale, $\mu_{\text{crit}}$, for the channel ${n_{\text{P}}+1}$ which is evanescent for energies below $\mu_{\text{crit}}$ and planar beyond $\mu_{\text{crit}}$. Consequently, momenta can no longer be approximated to be constant ${k_{n}\left(E\right)\not\approx\mathrm{const.}}$ if ${E\approx\mu_{n}}$. Instead, assuming the dispersion relation given in Eqs. (4) and (5) indicates that momenta ${k_{n}\left(E\right)}$ around the local extremum $\mu_{\text{crit}}$ are highly energy sensitive, following a square root behaviour. Note that this is true for energies both above and below $\mu_{\text{crit}}$. The only difference is that the channel ${n_{\text{P}}+1}$ either exhibits a strongly varying imaginary wave vector $\mathrm{i}\kappa_{n_{\text{P}}+1}(E)$ (in the evanescent regime) or a strongly varying real wave vector $k_{n_{\text{P}}+1}(E)$ (planar). In either case, taking the usual differentiability conditions, momentum drops down ${\sim k_{n}\left[E\right]\cdot\psi}$ and enters in first power. Accordingly, it is the momentum $k_{n_{\mathrm{P}}+1}\left(E\right)$ of the channel ${n_{\text{P}}+1}$  and this parameter only  by which energy dependence enters the scattering problem. If scattering is not translational invariant, as discussed above, the dependence of $k_{n_{\mathrm{P}}+1}\left(E\right)$ is projected onto all remaining channels via cross channel interaction, such that the overall energy dependence is dominated by the dispersion of the channel ${n_{\text{P}}+1}$. Thus, we can already predict on this general level, that the behaviour of the scattering matrix around that transition is given by the square root law ${\sim\sqrt{|1-E/\mu_{\text{crit}}|}}$.
@@
As a consequence, if the energy is close to the point where additional channels enter the scattering problem, the scattering coefficients smoothly, but strongly, adjust as a function of energy, with the square root power law. That is, there is a region of transition in the vicinity of $\mu_{\text{crit}}$ where the scattering matrix exhibits a strong algebraic energy dependence even below the point where the next higher channel joins the scattering process (cf. the gray stripe in Figure 5). The full picture, then, is that by approaching $\mu_{\text{crit}}$ from below, evanescent modes gradually become increasingly extended within the conductor arms and eventually turn into planar channels as the next higher mode is populated. In current treatments, it is assumed that evanescent modes exciting the scattering centre quickly fade away and thus can safely be neglected. However, as we show below in more detail, for short junctions this assumption is in general valid only for energies sufficiently far from $\mu_{\text{crit}}$. If the chemical potential is tuned to values close to a channel number transition, especially the lowest evanescent mode (whose spatial extension is diverging here) has to be taken into account, as it essentially pierces the SN-interfaces (cf. Figure 1). Crucially, this requires a generalisation of the interference condition, Eq. (17).
@@
Finally, the above analysis is not only important for the fundamental understanding of weak link physics in the general multi-channel case, but also useful for explicit calculations. As we show below with the concrete example, the square root behaviour can be used as an efficient way to interpolate the scattering matrix at different values of $E$, and thus significantly increase the computational speed.
III.5 Challenges for the generalisation of the gap closing argument
@@
As we just demonstrated, when the chemical potential is tuned close to a value where the number of planar channels changes (described by the small energy scale $\mu_{\text{crit}}$), we have a finite energy dependence of the scattering matrix, in spite of negligible dwelling time. It would therefore seem reasonable to expect that the discussion of Sec. III.3 applies here as well, in that the finite energy dependence of the scattering matrix would initially open a gap (i.e., detach $E$ from $\Delta$), but that a correction in $R$ would eventually restore the gap closing. And while ultimately, at least parts of that expectation turn out to be true, the pathway towards a full understanding of the near-gap spectral properties of the ABS in this regime is much more complicated.
@@
First, there is a simple energy scaling argument indicating that we need to explicitly include evanescent modes into the Beenakker framework. Namely, for the single channel case (see Sec. III.3) we saw that the only relevant energy scale is $\mu$, and that the energy-dependant corrections of the scattering matrix $\sim 1/\mu$ exactly cancelled with the dispersive correction in $R$ (likewise $\sim 1/\mu$). If we were to simply apply the same principle to the multi-channel case, but $\mu$ still tuned to values where there is only one planar channel, we now understand that the planar part of the scattering matrix has corrections that scale with ${1/\mu_{\text{crit}}\gg 1/\mu}$, and therefore, cannot simply be corrected by a mere dispersive correction to the planar Andreev reflection scaling with $1/\mu$. Indeed, as we show below, we need to take into account evanescent modes not only within the scattering matrix describing the central scattering region, but also the evanescent equivalent of Andreev reflection (that is, not only $S$ but also $R$ needs to be extended). Indeed, for the latter, the energy scale $\mu_{\text{crit}}$ emerges again, such that we have at our disposition a process depending on the same, dominant energy scale.
@@
This, however, would not be sufficient, as there is an even more profound challenge. Namely, with Eq. (29), we derived for the single channel case a constraint on the scattering matrix beyond unitarity (but still rooted in probability conservation), under the condition that the finite energy dependence in $S$ originated from a finite dispersion of the conduction electrons. Indeed, it was precisely this constraint that allowed us to show that $E=\Delta$ remained the correct solution at $\phi=0$ [see discussion around Eq. (31)]. It turns out that we can generalise the argument leading to Eq. (29) to the present multichannel case – but this generalised condition will not be helpful. Namely, one can extend the scattering wave function ansatz of Eqs. (68) and (69) to include the dominant evanescent mode in addition to the planar mode, with an evanescent version of a transition and reflection amplitude. However, note that in order for the scattering wave function to be normalisable, we can only include the decaying mode (and not the diverging mode) – an important caveat, as we will see in a moment. Now, proceeding in the same way as in Sec. III.2, one can derive a very similar type of condition as in Eqs. (78) and (81), except that the terms $\sim(R-R^{*})/\mu$ and ${\sim(T-T^{*})/\mu}$ would be replaced by terms proportional to the evanescent reflection and transmission amplitudes, divided by $\mu_{\text{crit}}$. Interestingly, this new constraint would tell a story in between the previously encountered situations. Here, the dip in the probability of the planar mode occupation upon impact of the wave packet at the scattering region is very much real. However, the total occupation probability of the electrons inside the conductor arms is still conserved, because at impact, the evanescent mode is temporarily occupied [cf. Figure 3b)].
@@
Yet, crucially, this constraint would not be a suitable condition to understand the properties of the ABS spectrum near the gap, for the simple reason that the ansatz for the full ABS wave function requires (contrary to the scattering wave function) both decaying as well as diverging parts of the evanescent modes (again, for details see below). The updated probability conservation constraint involving decaying modes however would only provide a relationship between the energy-dependence of the scattering amplitudes of the planar modes and the decaying evanescent modes, and thus provides only incomplete information. As a last resort, one could try to come up with an even more general ansatz for an incident wave packet including both decaying and diverging evanescent modes. But, it is impossible to construct normalisable wave packets in this way.
@@
By carefully generalising the Beenakker formalism to include evanescent modes below, we will be able to attack this problem from a different angle. In fact, in a sense, we will invert the problem by first providing a perturbative expansion of the generalised Beenakker equation up to first order in $1/\mu_{\text{crit}}$. Then, instead of providing a condition for the scattering matrix such that a gap closing at $\Delta$ is guaranteed, we demand that the gap closing exists, and deduce from that condition a constraint on the generalised scattering matrix (including evanescent modes). Crucially (as indicated above), this constraint cannot follow from probability conservation arguments. Thus, we provide an example where Andreev physics can provide us with fundamental information about the central scattering region, that cannot be obtained otherwise.
@@ IV Beenakker equation including evanescent modes
In order to include evanescent modes into the ansatz of the ABS wave function, we have to find generalisations for both $S$ and $R$ in Eqs. (1) and (10). As already indicated, we need to include in general both decaying and diverging evanescent modes within the unproximitised part of the conductor arms.
Overall, the wave function now needs to have a generalised notion of incoming and outgoing wave amplitudes, $\psi^{\text{in}}$ and $\psi^{\text{out}}$. This distinction is obvious only for planar modes, where one simply distinguishes the propagation direction via the group velocity. For evanescent modes, such a distinction is meaningless, such that we simply choose to define the solutions that decay away from the scattering matrix as outgoing, or “scattered”, evanescent modes (i.e., their amplitudes are added in $\psi^{\text{out}}$), and those that diverge as incoming, or “incident”, modes ($\psi^{\text{in}}$), see Fig. 6. We stress that this choice is arbitrary. The resulting wave function ansatz is nonetheless correct, provided that the scattering at the central scattering region, and at the SN interface, are computed consistently with the ensuing asymptotic boundary conditions away from the scattering region.
As a first step, the structure of the scattering matrix as shown in Eq. (6) has to be generalised to include evanescent modes. That is, instead of the submatrices $S^{n,m}$ being limited to indices $n,m$ going from $1$ to $n_{\mathrm{P}}$, where $n_{\mathrm{P}}$ is the number of planar modes, $n,m$ now have (at least in general) no upper bound. For illustration purposes, we explicitly separate the total (infinitely dimensional) scattering matrix into subblocks
@@
As just indicated, since there is an infinite number of higher evanescent modes, the S-matrix now becomes infinitely dimensional. Consequently, for explicit calculations, one has to find an appropriate cutoff for the number of modes, to have convergent low-energy solutions. As we see in a moment, for realistic problems, it actually often suffices to include only the first evanescent mode, whereas all higher modes can be neglected.
@@
With evanescent modes included into the model, there now are real exponentials exiting the scattering centre within the conductor arms. Note that conductor arms are often included as fictitious entities, and their size is ultimately set to zero [cf. Fig. 1b)]. However, some experimental device designs have structures that in principle allow for finite conductor arms, see Refs. [74, 44, 45], such that it is meaningful to also consider long junctions as in Fig. 1a). At any rate, as soon as evanescent modes pierce the proximitised S-region, we in general have to explicitly solve also the evanescent mode version of the Andreev scattering problem at the SN interface [^5] Since for $S$, we classified modes decaying from the scattering centre as “outgoing” (appearing in $\psi^{\text{out}}$) and diverging modes as “incoming” (appearing in $\psi^{\text{in}}$), we have to apply the same consistent grouping for the boundary conditions at the SN interface as well – now for both electrons and holes. While the wave function ansatz on the N side has both decaying and diverging evanescent components, on the S side, we again have to restrict the ansatz to only decaying modes when receding from the SN interface within the S region (see Fig. 6), to ensure normalisability of the total ABS wave function.. Moreover, in the standard framework, all coefficients were computed in the limit of a strongly linear dispersion relation (Andreev approximation), such that there was no electron-electron (or hole-hole) backscattering, and the Andreev reflection coefficient $\alpha$ did not depend on the channel index. All of these features must now be included. That is, instead of the form given in the standard framework by Eq. (11), the total reflection matrix at the SN interface now assumes the form
@@
The matrices $\widehat{\alpha}_{\nu}$ and $\widehat{\beta}_{\nu}$ are diagonal matrices whose size is given by the number of channels times the number of terminals, in analogy to the construction of the matrix ${\text{e}^{-\text{i}\widehat{\phi}}}$ as defined in Eq. (16). The matrix $\widehat{\alpha}_{\nu}$ generalises Andreev reflection, whereas $\widehat{\beta}_{\nu}$ captures the aforementioned normal backscattering. The coefficients on the diagonal depend on the channel index (but not the terminal index) due to the non-linearity of each channel’s dispersion. As we discuss in detail, there is in particular a marked difference in the behaviour of the coefficients for planar- and evanescent modes. As already discussed in Sec. III.3, finite normal backscattering (${\widehat{\beta}\neq 0}$) is important especially if it is the planar channels that exhibit strong dispersion. In what follows, we show that if dispersion instead stems from an evanescent mode, corrections to $\widehat{\alpha}$ will play a pivotal role.
@@
$\alpha_{\nu,n}$ $=2\nu\frac{\Delta}{E}\frac{K_{\nu,n}k_{\mathrm{R},n}}{k_{\mathrm{R},n}\zeta_{n}^{-}+\frac{\Delta}{E}\sqrt{1-\frac{E^{2}}{\Delta^{2}}}\Gamma_{n}^{\nu,+}}\ ,$ (48)
@@
$\Gamma_{n}^{\nu,\pm}=$ $\left[\left(k_{\mathrm{I},n}\pm K_{\nu,n}\right)\left(k_{\mathrm{I},n}+K_{-\nu,n}\right)+k_{\mathrm{R},n}^{2}\right]\text{ ,}$
$\zeta_{n}^{\pm}=$ $\left(K_{\mathrm{e},n}\pm K_{\mathrm{h},n}\right)\ .$ (50)
@@
The planar and evanescent momenta ${k_{\mathrm{e},n}}$ and ${k_{\mathrm{h},n}}$ are obtained by considering the dispersion in the N-region [cf. Eq. (4)] and solve it for the momentum $k_{n}$, where evanescent channels require the replacement ${k_{n}\to-\mathrm{i}\kappa_{n}}$, such that
@@
$\kappa_{\nu,n}$ $=\sqrt{2m\left(\epsilon_{n}-\mu-\nu E\right)}\ .$ (52)